OpenStudy (anonymous):

h(p(x)) h(x)= Sqrt(x+2) p(x)= x^2-2

OpenStudy (***[isuru]***):

first substitute p(x) instead of x in h(x) \[h[p(x)] = \sqrt{p(x) +2}\]

OpenStudy (***[isuru]***):

now.. substitute the value of p(x) in the above equation... and u will get ur answer :) can u do that ?

OpenStudy (anonymous):

Yes, I understand that part, but i need help solving it from there(:

OpenStudy (anonymous):

@***[ISURU]***

OpenStudy (***[isuru]***):

\[\sqrt{x^{2} - 2 + 2}\]\[\sqrt{x^{2}}\]remember this one ... \[\sqrt{a} = a^{\frac{ 1 }{ 2 } }\] so ..this will be \[x^{2 \times \frac{ 1 }{ 2 } } = x\] hope this will help ya!!